Find the resultant if the following forces acts on a body


12N along x-axis

8N along y-axis

4N at 30° along x-axis


1
Expert's answer
2021-05-04T12:07:49-0400

Find the components:


"F_{1x}=12\text{ N},\\
F_{1y}=0,\\
F_{2x}=0,\\
F_{2y}=8\text{ N},\\
F_{3x}=4\cos30°=3.46\text{ N},\\
F_{3y}=4\sin30°=2\text{ N}.\\"

The resultant:


"R_x=F_{1x}+F_{2x}+F_{3x}=15.46\text{ N},\\
R_y=F_{1y}+F_{2y}+F_{3y}=10\text{ N},\\
R=\sqrt{R_x^2+R_y^2}=18.41\text{ N},\\\space\\
\theta_R=\arctan\frac{R_y}{R_x}=32.9°."